| x+2 | =9

Can you show steps how to do this and then ill understnad how to do the rest of the questions. Thank you.

Answer by addingup(3677)

|x+2| = 9
Split into its two possible cases:
(x+2) = 9 or -(x+2) = 9
x+2 = 9 or -x-2 = 9
x = 7 or x = -11
Then, your answer is:
x = -11, 7